For Adam's tenth birthday, his grandparents opened a savings account in his name. They made an initial deposit of $750. During the year, Adam made no additional deposits and no withdrawals. At the end of the first year, Adam's money had earned $26.25. Use the formula I = Prt, where I is the amount of interest earned, P is the principal, or the original amount deposited, r is the interest rate, and t is the time for which the money is invested to determine what the interest rate is on his savings account.
I know the equations but forgot the steps, can you walk me through it?
Here is my work so far: $26.25= 750r12
I used 12 in order to replace the 365 days ( 12 months)
What do I do now since the r is in the middle?

Answer by stanbon(75887)

For Adam's tenth birthday, his grandparents opened a savings account in his name. They made an initial deposit of $750. During the year, Adam made no additional deposits and no withdrawals. At the end of the first year, Adam's money had earned $26.25. Use the formula I = Prt, where I is the amount of interest earned, P is the principal, or the original amount deposited, r is the interest rate, and t is the time for which the money is invested to determine what the interest rate is on his savings account.
----
I = P*r*t
---
Note: The "r" is an annual rate
The t is stated in years.
--------------
26.25 = 750*r*1
-----
r = 26.25/750
-----
rate = 0.035 = 3.5%
=====================
Cheers,
Stan H.
====================